OpenStudy (anonymous):

What is the equation of a line that passes through the point (-3, 1) and is parallel to a line with a slope of 2?

OpenStudy (sheraz12345):

Apply formula y – y1 = m(x – x1) Slope will be 2 as lines are parallel

OpenStudy (anonymous):

y = 2x + 1 y = 2x + 4 y = 2x + 5 y = 2x + 7

OpenStudy (anonymous):

need more help

OpenStudy (sheraz12345):

Third opt

OpenStudy (sheraz12345):

Did u get it @edjoey

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

where did u get the 5 from?

OpenStudy (sheraz12345):